Operation Summary for December 1, 2010 - December 3, 2010 |
Wednesday Dec. 1 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (355.7E10), with the Antiproton Source (Pbar) stacking (31.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations established HEP store 8321 at 8:40 AM with an average initial luminosity (AIL) of 321E30. Operations established beam to the next Meson FTBF experiment, T-1008 at 9:45 AM. Operations reported that a Linac RF station (LRF1) tripped off due to a failed fuse in its driver anode cabinet. During the evening shift, Operations disabled beam to T-1008 at 6:15 PM. Operations reported that a programmable logic controller (PLC) for NuMI failed and returned incorrect LCW data for the tank level. Controls personnel reprogrammed the PLC to look at the backup system. Store 8321 was not terminated at the scheduled time due to a failed pond water pump for the D0 experiment. D0 personnel turned off their detector to cut the heat load. FESS personnel investigated and determined that the inlet water pipe was clogged. Thursday Dec. 2 The midnight shift began with Operations monitoring store 8321 (72.55E30) and the stash (378.2E10), with Pbar stacking (12.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 1:30 AM, D0 personnel indicated the investigation into the clogged pond water inlet pipe would continue throughout the shift. Operations terminated store 8321 at 1:50 AM, put the TeV into shot setup, and established store 8322 at 3:55 AM with an AIL of 307E30. At 6:25 AM, Operations reported that the Linac critical devices closed due to bad beam positions. Operators investigated and determined that the problem was with the 400 MeV area chopper. Operations contacted Booster personnel, who continued the investigation. During the day shift, Booster personnel re-terminated the 400 MeV chopper On-Tube cable. The Proton Source resumed delivering beam at 8:35 AM. During the evening shift, Operations disabled beam to T-1008 at 4:35 PM. D0 personnel calibrated their detector. Operations put the TeV into shot setup at 6 PM and established store 8323 at 7 PM with an AIL of 319E30. Operators rebooted the MI ramp control crate (MECAR) to resolve some memory issues. All MI ramps had to be re-learned into the memory. Friday Dec. 3 The midnight shift began with Operations monitoring store 8323 (142.7E30) and the stash (152.7E10), with Pbar stacking (24.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that a Switchyard (SY) vacuum roughing pump (110RP) had tripped off and would not turn back on. Operators contacted vacuum technicians and, with the technicians help, determined the cause to be a faulty status and control card; Operators replaced the card at 6:35 AM. Operations then established beam to T-1008. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 12/03/2010 email Fermilab |
| Security, Privacy, Legal | |